OpenStudy (anonymous):

Simplify the product. (x-4)(x+3) A.x^2-7x-12 B. x^2+x-12 C. x^2-x-12 D. x^2-12x-1 HELP PLEASE!

OpenStudy (anonymous):

(x-4)(x+3) Can you use FOIL to expand them?

OpenStudy (anonymous):

it is usually C, and in fact this time it is C

OpenStudy (anonymous):

\[(x-4)(x+3)=x\times x+3\times x-4\times x-4\times 3=x^2-x-12\]